how many atoms of carbon are there in 0.40 mol of procaine, C13H20N2O2, a "pain killer" used by dentist?
1
Expert's answer
2016-02-20T00:00:58-0500
N(C13H20N2O2) = c* Na = 0.4 * 6.02*1023 = 2.4 *10^23
N(C) = N(C13H20N2O2) * 13 = 3.13*10^24
Answer
N(C) = 3.13*10^24